OpenStudy (anonymous):

Which system has (-8, 0) as the solution? Which system has infinite solutions? Which system has (1, 3) as the solution? Which system has no solutions? Which system has (-2, -2) as the solution?

OpenStudy (anonymous):

A) x + y = -4 -3x + 2y = 2 B) x - 4y = -8 2x - 3y = -16 C) x - y = -2 3x - y = 0 D) x + y = -1 4x + 4y = -4 E) 4x + 6y = -12 2x + 3y = 9

OpenStudy (saifoo.khan):

A = Which system has (-2, -2) as the solution?

OpenStudy (saifoo.khan):

B = Which system has (-8, 0) as the solution?

OpenStudy (saifoo.khan):

C = Which system has (1, 3) as the solution?

OpenStudy (saifoo.khan):

D Which system has infinite solutions?

OpenStudy (saifoo.khan):

E = No solution